John is 12, Peter is 3. How many years will it take for John to be 3 times as old as
Peter?

Answer by ikleyn(52786)

.

In x years.


In x years, John will be 12+x years old;  Peter will be 3+x years old.


John will be 3 times as old as Peter:


    12 + x = 3*(3 + x).


Simplify and find x


    12 + x = 9 + 3x

    12 - 9 = 3x - x

       3   = 2x

       x    = 3/2 years.


When the age problem has no solution in integer numbers, it is TRADITIONALLY considered as DEFECTIVE.

Alas (!)

-------------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - HOW TO algebreze and to solve age problems?
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Age problem for three participants
    - Miscellaneous age problems
    - Entertainment age problems
in this site.

Read them, have fun and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.